OpenStudy (anonymous):

sonya has 10 dollars less then paula. together they have 90 dollars. would the equation be s = p -10 = 90 and how would you solve it? step by step plz :)

OpenStudy (agreene):

(p-10)+p = 90 would be a better interpretation of the information. p-10+p=90 combine likes 2p-10=90 add 10 to both sides 2p=100 divide by 10 p=50

OpenStudy (anonymous):

here one change sonya has x dollars so paula has x+10 dollars now u can solve it!!

OpenStudy (anonymous):

did u got wht mistake u were makin

OpenStudy (anonymous):

yess thank u